OpenStudy (anonymous):

Write the solution set for equation 7x - 2y = 8 with this replacement set: {(1,-1),(2,3), (0,-4), (4,10)}. List your answers from least to greatest based on the x value and without any spaces.

OpenStudy (owlcoffee):

Any point in the cartesian coordinates actually has the structure: \[A(x , f(x))\] But, since it's unnecesarly complex to actually look at it like that, let's say that: \[y=f(x)\] therefore: \[A(x,y)\] So, moving on to the problem in question, we have the equation of: \[7x-2y=8\] Now, something very very important when we evaluate points i a given equality, and that is that in order for it to be a solution it has to satisfy the equation. What do I mean? By that I mean. that when I replace the point, it has to give me the constant that it it actually equal to. Lt's say for example, a very general equation: \[Ax+By=C\] And we say that point P(m,n) acutally is a solution for that equation, then the following must be true: \[A(m)+B(n)=C\] The only thing i did was replace the coordinates of the point in te equation, and for it to be a solution, the number it gives me must be equal to the C on the left side of the equality. A, B, C, m, n are any real number you like. So, let's return: \[7x-2y=8\] let's take one of the numbers we are given and replace it there, say (1,-1). Now, I already told you, that any point has the structure of A(x,y), so by simply looking you can see that x=1 and y=-1. So, let's take the equation 7x-2y=8 and instead of x and y let's put 1 and -1 . It'll look like this: \[7(1)-2(-1)=8\] \[7+2=8\] \[9 \neq 8\] Since 9 is not equal to 8, (1,-1) is not a solution for the equality.